A mass m , free to move along a line is attracted towards a given point on the line with a force proportional to its distance from the given point.If the mass starts from rest at a distance x_o from the given point, show that the mass moves in a simple harmonic motion?
1
Expert's answer
2019-09-26T11:14:29-0400

Introduce a Cartesian coordinate system, where the center of attraction is origin and "\vec{r}(0)=x_0\vec{e}_1"

We have "m\vec{r}''=-\alpha\vec{r}" for some "\alpha>0" . Let "\frac{\alpha}{m}=\omega^2" , then "\vec{r}''+\omega^2\vec{r}=0".

If "\vec{r}=(x_1,\ldots,x_n)", then "x_i''+\omega^2 x_i=0" for all "i=1,\ldots,n" .

So for all "i=1,\ldots,n" we have the general solution "x_i=B_i\cos\omega t+C_i\sin\omega t="

"\frac{1}{\sqrt{B_i^2+C_i^2}}\bigl(\frac{B_i}{\sqrt{B_i^2+C_i^2}}\cos\omega t+\frac{C_i}{\sqrt{B_i^2+C_i^2}}\sin\omega t\bigr)="

"=A_i\cos(\omega t+\omega_i)" , where "A_i=\frac{1}{\sqrt{B_i^2+C_i^2}}" and "\omega_i" such number that "\cos\omega_i=\frac{B_i}{\sqrt{B_i^2+C_i^2}}" and "\sin\omega_i=-\frac{C_i}{\sqrt{B_i^2+C_i^2}}"

We obtain "\vec{r}(t)=(A_1\cos(\omega t+\omega_1),\ldots,A_n\cos(\omega t+\omega_n))".

Initial condition gives us "(x_0,0,\ldots,0)=\vec{r}(0)=(A_1\cos\omega_1,\ldots,A_n\cos\omega_n)", so "A_1\cos\omega_1=x_0" and "A_i\cos\omega_i=0" while "i>1".

This means that 1)"A_1\cos\omega_1=x_0" and 2)"A_i=0" or "\cos\omega_i=0" for all "i>1".

Anyway, it was obtained that "\vec{r}(t)=(A_1\cos(\omega t+\omega_1),\ldots,A_n\cos(\omega t+\omega_n))" is an equation of simple harmonic motion.